Operations Report – December 2nd

Crew 216 Operations Report 02-Dec-19
SOL: 8
Name of person filing report: Mike Lawson

Non-nominal systems: NA
Generator: run
Hours run: 12.25
From what time last night: 1845
To what time this morning: 0700
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 52 %

Diesel Reading – 36
Station Propane Reading – 54
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1464017
Water (static tank) (gal): 269
Static to Loft Pump used – yes
Water in Green Hab (gal): 255
Water in Science Dome (gal): 0

Toilet tank emptied: no

Deimos rover used: Off Campus for repairs
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: No
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: yes
Hours: 115.7
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 38%
Currently charging: yes

Opportunity rover used: Off Campus for repairs
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging:

Curiosity rover used: yes
Hours: 120.7
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 35.%
Currently charging: yes

Notes on rovers: Opportunity to be moved off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where?

General notes and comments: EVA North Ridge, radio relay site surveyed. EVA to vicinity quarry.

Summary of internet: No serious issues.

Summary of suits and radios: No issues.

Summary of Hab operations: Routine maintenance and housekeeping tasks. Replacement of hot water tank installed, running in 120 VAC mode only. Needs an additional switch to run on propane. Will not be an issue when the station is on the generator. Generator frequency is slightly off, electric clocks run fast by 10-15 minutes/day. Not an issue when on the battery bank.

Summary of GreenHab operations: routine watering.

Summary of Science Dome operations: NA

Summary of RAM operations: RAM Rover power charging stations inoperable, electrical issues remain unresolved. Director is aware of issue.

Summary of any observatory issues: NA

Summary of health and safety issues: NA

Questions, concerns and requests to Mission Support: 2 bag NB, 1 bag B trash in the rear airlock.

Copyright © The Mars Society. All rights reserved. | Main Site